Show MetadataProvide my evaluation of this resource
Which of the following statements about the ordinary IR spectroscopic regions are TRUE?
True False  In general, the energy carried by a photon of IR radiation is greater than radiation is greater than the energy carried by a photon of UV radiation.
True False  Ordinarily we sense most IR frequencies as colors in the region we call yellow.
True False  We can sense some of the frequencies of the FUNDAMENTAL region of the IR as heat.
True False  In general, the IR FUNDAMENTAL region has a longer wavelength region than the region we call the ultraviolet (uv) region.
True False  An IR wavelength of 40 microns corresponds to a wavenumber of 2500 inverse cm.
Answer #1

In general, the energy carried out by photon of IR Radiation is less than energy carried out by photon of UV radiation.So,the wavelength of IR radiation has longer wavelength than wavelength of UV radiation.It can be understood by following relation.

E = hc

wavenumber=E=1/1   =1/40*10^(-4)=250 inverse cm

(IR Frequencies are colors in the region we call yellow) False statement

We can sense some of the frequencies of the FUNDAMENTAL region of the IR as heat. this is a true statement because it lowers energy consumption.

Thus, Statement I False, statement II False, Statement III True, Statement IV True,Statement V False